OpenStudy (shesolitt):

I WILL MEDAL AND FAN Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2. −14 −7 7 11

OpenStudy (shesolitt):

@Vuriffy

OpenStudy (helpppp):

fill in for x

OpenStudy (helpppp):

-4-3=-7 -2+4=2 2(7)

OpenStudy (helpppp):

2(-7) ffs

OpenStudy (shesolitt):

i got 11 @helpppp

OpenStudy (helpppp):

2(-7) is -14

OpenStudy (helpppp):

solve for g(x) first

OpenStudy (helpppp):

-2(2) is -4 -4-3 is -7

OpenStudy (helpppp):

Now solve for f(x)

OpenStudy (helpppp):

-(2)+4 is 2

OpenStudy (helpppp):

mu;tiply the 2 solutions together

OpenStudy (vuriffy):

f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2. -(2) + 4 = 2. -2(2) - 3 = -7. 2(-7) = -14.

OpenStudy (will.h):

Need help?

OpenStudy (shesolitt):

so the answer is 14 @vu

OpenStudy (shesolitt):

@Vuriffy

OpenStudy (helpppp):

-14..... a positive times a negative is a negativeeee

OpenStudy (will.h):

You look nervous helpppp

OpenStudy (shesolitt):

im confused

OpenStudy (helpppp):

Im not nervous lol. I wish that pic was me

OpenStudy (shesolitt):

is -14 correct

OpenStudy (helpppp):

yesss

OpenStudy (helpppp):

2(-7) is -14 YOUR ANSWER IS A

OpenStudy (will.h):

I think what confuses me is that I have no clue if you guyes are solving 2 separate questions or one

OpenStudy (shesolitt):

thank you everyone !